OpenStudy (samigupta8):

let K be an integer and p is a prime such that the quadratic equation x^2+Kx+P=0 has two distinct positive integer solution.The value fo K-P equals to

OpenStudy (samigupta8):

@AllTehMaffs pls...hlp

OpenStudy (samigupta8):

i also did upto dat and afterward my brain stopped wrking

OpenStudy (anonymous):

\[K-P > 2 \sqrt P - P\] o_0

OpenStudy (anonymous):

but P is a prime, so there's no such integer as sqrt P

OpenStudy (samigupta8):

wt i m notr able to undrstand

OpenStudy (anonymous):

hmmm \[K^2 -4P > 0\] \[(K+2P)(K-2P)>0\] so \[K>-2P\] \[K>2P\] ?

OpenStudy (anonymous):

this isn't helpful I think :P

OpenStudy (anonymous):

if P is prime then only 1 divides it, so \sqrt P can't be an integer. But I don't think that's helpful in this problem.

OpenStudy (samigupta8):

the answer is 1 bt how

OpenStudy (anonymous):

Well the only solution that satisfies this inequailty \[K-P > 2\sqrt P - P\] is P = 1, then the answer on the right hand side jas to be 1

OpenStudy (samigupta8):

hw u did that

OpenStudy (anonymous):

given that k is an integer and p is a prime so......both the roots will be integers coz considering the case of fractional can satisfy the condn of K but not of P.....so we know that roots are integers.... then multiplying two integers can be only prime when one of them is 1 otherwise its not possible, so we have one root as 1

OpenStudy (anonymous):

then K= (1+a) and P = 1*a...... K-P = 1

OpenStudy (samigupta8):

that's certainly gud .......

OpenStudy (anonymous):

the way I did it was \[K^2 - 4P > 0\] because of the quadratic formula, then \[K^2 > 4P\] \[K>2\sqrt P\] \[K-P > 2\sqrt P - P\] The only prime that has an integer square root is 1, , so P=1 \[K-P > 2-1\] K-P > 1 But it's still an inequality; I think div.u might have a more solid solution

OpenStudy (primeralph):

Am I interrupting?

OpenStudy (samigupta8):

no bt i got da answer @divu.mkr explained me

OpenStudy (primeralph):

Okay then, good luck.